# VisualCloze: A Universal Image Generation Framework via Visual In-Context Learning

### v2

@@ 4.3 Implementation Details
We use FLUX.1-Fill-dev [33] as our foundation model, considering its outstanding performance among open-source image infilling models. In this work, LoRA [25] is chosen to fine-tune the model instead of fully fine-tuning it to reduce training costs and preserve the capabilities of the foundation model. The resulting LoRA can also be fused with other LoRAs in the community, enabling more widespread applications. Specifically, we set the rank of LoRA as 256256256. The model is tuned for 20,000 iterations with an accumulated batch size of 64 on 8 $\times$ A100 GPUs. We employ the AdamW optimizer with a learning rate of $1e^{-4}$. Following FLUX.1-Fill-dev, we incorporate the lognorm noise strategy with dynamic time shifting. During training, the number of in-context examples is set up to 2 (i.e., $C$ as defined in Sec. 4.2), while $L$, the number of images involved in a task, varies between 2 and 4 in the Graph200K dataset. During inference, the number of in-context examples can be generalized to a larger number. To balance computational efficiency, each image is resized to the area of $\times 384$ 384 or $\times 512$ 512 before concatenating them into a grid layout. High-resolution outputs can be obtained in practical applications through simple post-up-scaling techniques [45].
@@ 5.1 Qualitative Analysis of In-context Learning
Generalization on unseen tasks. Beyond mitigating task confusion, in-context learning also enables the model to generalize to tasks unseen during training. Fig. 2 has shown the model can successfully generate frontal faces from side-view images and transfer editing instructions [8] through in-context learning, even though they are not encountered during training. Here, we present additional examples of unseen tasks. For instance, although the model is trained exclusively on image editing tasks involving object addition and removal, it still generalizes to other types of editing tasks, such as environment changes and attribute modifications, as shown in Fig. 7. Furthermore, as demonstrated in Fig. 8, the model, trained solely on single-subject generation, can generate images preserving identities of multiple subjects. These results highlight that in-context learning is an effective guidance mechanism, enabling adaptation to novel tasks without retraining.
[…]
@@
Varying effects of different in-context examples. Following prior works [52, 46] on the prompt selection, we also find that different in-context examples could impact the generation quality. Specifically, it is crucial that in-context examples provide correct and strong guidance about the task intention. For example, as shown in Fig. 10 (left), when the side faces are more towards the front than in Fig. 10 (right), the success rate of correctly generating frontal faces has dropped dramatically.
@@
[…]
@@
[…]
@@
Canny ControlNet [80] 0.13 - 46.06 0.34 0.31 45.45 34.10
OminiControl [61] 0.47 - 29.58 0.61 0.44 61.40 34.40
OneDiffusion [35] 0.39 - 32.76 0.55 0.46 59.99 34.99
OmniGen [71] 0.43 - 51.58 0.47 0.47 62.66 33.66
@@
Depth ControlNet [80] - 23.70 36.83 0.41 0.44 60.17 34.49
OminiControl [61] - 21.44 36.23 0.52 0.44 60.18 34.08
OneDiffusion [35] - 10.35 39.03 0.49 0.49 60.49 34.71
OmniGen [71] - 15.07 86.08 0.26 0.49 64.90 29.72
@@
Deblur ControlNet [80] - 37.82 53.28 0.49 0.45 61.92 33.80
OminiControl [61] - 19.70 26.17 0.85 0.45 60.70 34.53
OneDiffusion [35] - - - - - - -
OmniGen [71] - - - - - - -
@@
OminiControl [61] 73.17 87.70 33.53
OneDiffusion [35] 73.88 86.91 34.85
OmniGen [71] 67.73 83.43 34.53
@@
   $\uparrow$    $\uparrow$
    InstantStyle [64]     0.27     0.60
   OmniGen [71]    0.27    0.52
@@ 5.2 Main Results
We compare our method with universal generative models, including OmniGen [71] and OneDiffusion [35], as well as specialized models, such as ControlNet [80] and OminiControl [61]. The details of the evaluation metrics are provided in Appendix C. Additionally, we fine-tune FLUX.1-dev [33] using the same settings as FLUX.1-Fill-dev for comparison and refer to the tuned models as Oursdev and Oursfill. The details of Oursdev are shown in Appendix B.
For conditional generation and image restoration, we evaluate the models based on three criteria, i.e., controllability, visual quality, and text consistency, following the evaluation approach of OminiControl [61]. As shown in Tab. 1, our framework demonstrates comparable controllability to existing universal methods while achieving superior visual quality and text consistency. Compared to specialized methods, our model performs on par with the best results and even outperforms them on the depth-to-image.
In the style transfer task, we measure text consistency and style alignment using the CLIP [49] model. As reported in Tab. 3, our method outperforms OmniGen [71] by 2% and 3% in text alignment and style consistency, respectively. Even when compared with InstantStyle-Plus [81], a specialized model, we achieve a 2% improvement in text consistency, with only a slight decrease in style alignment.
Furthermore, we evaluate the models on subject-driven image generation and report semantic alignment using the DINOv2 [47], CLIP-I [49], and CLIP-T [49] scores. Across all these metrics, our method consistently delivers improvements, as shown in Tab. 2. For example, compared to the specialized model OminiControl [61], we achieve improvements of 7.15%, 1.66%, and 1.48% in these three scores.
Advantages of the infilling model. Our method (Oursfill) is built on FLUX.1-Fill-dev [33], which shares the same objective as our unified image generation framework. To verify its effectiveness, we also fine-tune Fill.1-dev [33] (Oursdev) using identical settings. Unlike Oursfill, which requires no modifications, Oursdev necessitates model adaptations for universal image generation, as shown in Appendix B. Despite its simplicity, Oursfill achieves superior performance across multiple tasks.
@@ Appendix A Instruction Format
Apart from FLUX.1-Fill-dev, we also adapt our method to FLUX.1-dev [33], a common text-to-image generative model. Unlike the infilling model that shares a consistent objective with universal image generation, FLUX.1-dev requires customized modifications to process clean condition images and noise target images. Specifically, after concatenating images in a grid layout like the infilling model, we always keep the region corresponding to the conditions as clean latent embeddings throughout the sampling process. This strategy requires modifications in image sampling because FLUX.1-Fill-dev takes noise latent embeddings as input. Moreover, for the adaLN-Zero block [48], it is critical to calculate the separate mean and shift parameters for the regions of clean conditions and noise target by feeding $T=0$ and $T=t$ into the adaLN-Zero, respectively. $t$ indicates the timestep in each sampling step and gradually increases from 0 to 1 along the sampling process. This strategy aligns with the pre-training domain of FLUX.1-dev, where different noise levels correspond to different mean and shift. As shown in Fig. 14, this strategy ensures the visual fidelity.
@@ C.1 Conditioning Generation
Generation quality. We measure the Generation quality using FID [23], SSIM, MAN-IQA [75], and MAN-IQA [75]. FID [23] measures the similarity between generated and real image feature distributions. SSIM evaluates perceptual quality by comparing luminance, contrast, and structural patterns between images. It calculates local patch statistics and combines them into a composite score ranging from 1-1- 1 to 111, with higher values indicating better structural preservation. MANIQA [75] and MUSIQ [30] leverage neural networks to predict image quality scores.
@@
Leveraging the powerful multi-modal capability of CLIP [49], we also measure the semantic alignment between generated images and text prompts, which reflects how the model follows instructions.
@@ C.2 Subject Driven Generation
Following DreamBooth [54] and BLIP-Diffusion [36], we measure DINOv2 [47], CLIP-I [49], and CLIP-T scores for the comparison of subject-driven image generation. DINOv2 [47] and CLIP-I scores measure the alignment between the reference subject and generated images through cosine similarity and CLIP score, respectively. CLIP-T measures the alignment between the generated image and the corresponding text prompt.
@@ C.3 Style Transfer
Following StyleDrop [56], we assess the performance of style transfer according to text consistency and style alignment. For text alignment, we measure the cosine similarity between embeddings of generated images and text prompts, where the embeddings are extracted by CLIP [49]. Regarding style consistency, we measure the cosine similarity between embeddings of generated images and style reference. Note that these two metrics should be considered together because the style consistency will reach 1.0 if the model collapses, where the model completely copies style reference as a composite image and ignores text instructions.

### v3

@@ 4.3 Implementation Details
We use FLUX.1-Fill-dev [flux2024] as our foundation model, considering its outstanding performance among open-source image infilling models. In this work, LoRA [hu2022lora] is chosen to fine-tune the model instead of fully fine-tuning it to reduce training costs and preserve the capabilities of the foundation model. The resulting LoRA can also be fused with other LoRAs in the community, enabling more widespread applications. Specifically, we set the rank of LoRA as $256$. The model is tuned for 20,000 iterations with an accumulated batch size of 64 on 8 $\times$ A100 GPUs. We employ the AdamW optimizer with a learning rate of $1e^{-4}$. Following FLUX.1-Fill-dev, we incorporate the lognorm noise strategy with dynamic time shifting. During training, the number of in-context examples is set up to 2 (i.e., $C$ as defined in Sec. 4.2), while $L$, the number of images involved in a task, varies between 2 and 4 in the Graph200K dataset. During inference, the number of in-context examples can be generalized to a larger number. To balance computational efficiency, each image is resized to the area of $384\times 384$ or $512\times 512$ before concatenating them into a grid layout. High-resolution outputs can be obtained in practical applications through simple post-up-scaling techniques [meng2021sdedit].
@@ 5.1 Qualitative Analysis of In-context Learning
Figure 7: Unseen Tasks U:  Although the image editing tasks seen by the model are only about object addition and object removal, it can still generalize to other types of editing tasks, such as environment modification (Left) and attribute transformation (Right), through in-context learning. More unseen tasks are shown in Fig. 2.
Generalization on unseen tasks. Beyond mitigating task confusion, in-context learning also enables the model to generalize to tasks unseen during training. Fig. 2 has shown the model can successfully generate frontal faces from side-view images and transfer editing instructions [chen2025edittransferlearningimage] through in-context learning, even though they are not encountered during training. Here, we present additional examples of unseen tasks. For instance, although the model is trained exclusively on image editing tasks involving object addition and removal, it still generalizes to other types of editing tasks, such as environment changes and attribute modifications, as shown in Fig. 7. Furthermore, as demonstrated in Fig. 8, the model, trained solely on single-subject generation, can generate images preserving identities of multiple subjects. These results highlight that in-context learning is an effective guidance mechanism, enabling adaptation to novel tasks without retraining.
@@
Figure 8: Unseen Tasks U: VisualCloze is capable of performing multi-subject driven generation [wu2025lesstomoregeneralizationunlockingcontrollability], even though the model was only exposed to single subject-driven generation tasks during training. Best viewed by zooming in.
@@
Varying effects of different in-context examples. Following prior works [rubin2022learningretrievepromptsincontext, 10172590] on the prompt selection, we also find that different in-context examples could impact the generation quality. Specifically, it is crucial that in-context examples provide correct and strong guidance about the task intention. For example, as shown in Fig. 10 (left), when the side faces are more towards the front than in Fig. 10 (right), the success rate of correctly generating frontal faces has dropped dramatically.
@@
Figure 9: Unseen Tasks U: Through in-context learning, we can perform reverse generation from targets to conditions. For example, (a) decomposing the layout and style from a stylized image and (b) inferring the image, depth, and surface normal simultaneously from an edge map, which is the reverse task of Fig. 3 (Left).
@@
Figure 11: Unseen Tasks U: Unseen combinations of multiple tasks. For conditional generation, we integrate multiple conditions achieve more precise control. More examples are shown in Fig. 3.
@@
Canny ControlNet [zhang2023adding] 0.13 - 46.06 0.34 0.31 45.45 34.10
OminiControl [tan2024ominicontrol] 0.47 - 29.58 0.61 0.44 61.40 34.40
OneDiffusion [le2024diffusiongenerate] 0.39 - 32.76 0.55 0.46 59.99 34.99
OmniGen [xiao2024omnigen] 0.43 - 51.58 0.47 0.47 62.66 33.66
@@
Depth ControlNet [zhang2023adding] - 23.70 36.83 0.41 0.44 60.17 34.49
OminiControl [tan2024ominicontrol] - 21.44 36.23 0.52 0.44 60.18 34.08
OneDiffusion [le2024diffusiongenerate] - 10.35 39.03 0.49 0.49 60.49 34.71
OmniGen [xiao2024omnigen] - 15.07 86.08 0.26 0.49 64.90 29.72
@@
Deblur ControlNet [zhang2023adding] - 37.82 53.28 0.49 0.45 61.92 33.80
OminiControl [tan2024ominicontrol] - 19.70 26.17 0.85 0.45 60.70 34.53
OneDiffusion [le2024diffusiongenerate] - - - - - - -
OmniGen [xiao2024omnigen] - - - - - - -
@@
OminiControl [tan2024ominicontrol] 73.17 87.70 33.53
OneDiffusion [le2024diffusiongenerate] 73.88 86.91 34.85
OmniGen [xiao2024omnigen] 67.73 83.43 34.53
@@
[…]
    InstantStyle [wang2024instantstyle]     0.27     0.60
   OmniGen [xiao2024omnigen]    0.27    0.52
@@ 5.2 Main Results
We compare our method with universal generative models, including OmniGen [xiao2024omnigen] and OneDiffusion [le2024diffusiongenerate], as well as specialized models, such as ControlNet [zhang2023adding] and OminiControl [tan2024ominicontrol]. The details of the evaluation metrics are provided in Appendix C. Additionally, we fine-tune FLUX.1-dev [flux2024] using the same settings as FLUX.1-Fill-dev for comparison and refer to the tuned models as Oursdev and Oursfill. The details of Oursdev are shown in Appendix B.
For conditional generation and image restoration, we evaluate the models based on three criteria, i.e., controllability, visual quality, and text consistency, following the evaluation approach of OminiControl [tan2024ominicontrol]. As shown in Tab. 1, our framework demonstrates comparable controllability to existing universal methods while achieving superior visual quality and text consistency. Compared to specialized methods, our model performs on par with the best results and even outperforms them on the depth-to-image.
In the style transfer task, we measure text consistency and style alignment using the CLIP [radford2021learning] model. As reported in Tab. 3, our method outperforms OmniGen [xiao2024omnigen] by 2% and 3% in text alignment and style consistency, respectively. Even when compared with InstantStyle-Plus [Zhang_2023_inst], a specialized model, we achieve a 2% improvement in text consistency, with only a slight decrease in style alignment.
Furthermore, we evaluate the models on subject-driven image generation and report semantic alignment using the DINOv2 [oquab2023dinov2], CLIP-I [radford2021learning], and CLIP-T [radford2021learning] scores. Across all these metrics, our method consistently delivers improvements, as shown in Tab. 2. For example, compared to the specialized model OminiControl [tan2024ominicontrol], we achieve improvements of 7.15%, 1.66%, and 1.48% in these three scores.
Advantages of the infilling model. Our method (Oursfill) is built on FLUX.1-Fill-dev [flux2024], which shares the same objective as our unified image generation framework. To verify its effectiveness, we also fine-tune Fill.1-dev [flux2024] (Oursdev) using identical settings. Unlike Oursfill, which requires no modifications, Oursdev necessitates model adaptations for universal image generation, as shown in Appendix B. Despite its simplicity, Oursfill achieves superior performance across multiple tasks.
@@ 7 Conclusion
[…]
Figure 13: Examples of language instructions that contain prompts about the layout of the concatenated image, task intent, and content of the target image.
[…]
Figure 14: Effects of separate mean and shift in fine-tuning FLUX.1-dev.
@@ Appendix A Instruction Format
Apart from FLUX.1-Fill-dev, we also adapt our method to FLUX.1-dev [flux2024], a common text-to-image generative model. Unlike the infilling model that shares a consistent objective with universal image generation, FLUX.1-dev requires customized modifications to process clean condition images and noise target images. Specifically, after concatenating images in a grid layout like the infilling model, we always keep the region corresponding to the conditions as clean latent embeddings throughout the sampling process. This strategy requires modifications in image sampling because FLUX.1-Fill-dev takes noise latent embeddings as input. Moreover, for the adaLN-Zero block [Peebles_2023_ICCV], it is critical to calculate the separate mean and shift parameters for the regions of clean conditions and noise target by feeding $T=0$ and $T=t$ into the adaLN-Zero, respectively. $t$ indicates the timestep in each sampling step and gradually increases from 0 to 1 along the sampling process. This strategy aligns with the pre-training domain of FLUX.1-dev, where different noise levels correspond to different mean and shift. As shown in Fig. 14, this strategy ensures the visual fidelity.
@@ C.1 Conditioning Generation
Generation quality. We measure the Generation quality using FID [heusel2017gans], SSIM, MAN-IQA [yang2022maniqa], and MAN-IQA [yang2022maniqa]. FID [heusel2017gans] measures the similarity between generated and real image feature distributions. SSIM evaluates perceptual quality by comparing luminance, contrast, and structural patterns between images. It calculates local patch statistics and combines them into a composite score ranging from 1-1 to $1$, with higher values indicating better structural preservation. MANIQA [yang2022maniqa] and MUSIQ [ke2021musiq] leverage neural networks to predict image quality scores.
@@
Leveraging the powerful multi-modal capability of CLIP [radford2021learning], we also measure the semantic alignment between generated images and text prompts, which reflects how the model follows instructions.
@@ C.2 Subject Driven Generation
Following DreamBooth [ruiz2023dreambooth] and BLIP-Diffusion [li2023blip], we measure DINOv2 [oquab2023dinov2], CLIP-I [radford2021learning], and CLIP-T scores for the comparison of subject-driven image generation. DINOv2 [oquab2023dinov2] and CLIP-I scores measure the alignment between the reference subject and generated images through cosine similarity and CLIP score, respectively. CLIP-T measures the alignment between the generated image and the corresponding text prompt.
@@ C.3 Style Transfer
Following StyleDrop [sohn2023styledrop], we assess the performance of style transfer according to text consistency and style alignment. For text alignment, we measure the cosine similarity between embeddings of generated images and text prompts, where the embeddings are extracted by CLIP [radford2021learning]. Regarding style consistency, we measure the cosine similarity between embeddings of generated images and style reference. Note that these two metrics should be considered together because the style consistency will reach 1.0 if the model collapses, where the model completely copies style reference as a composite image and ignores text instructions.